Suppose N=x^+2 models the number ofcases of infection, in millions, of a disease x years from now. Approximately how many years will there be 7 million cases.
I got this 7=x^+2
(-x^2)=(-5)
x=25

Answer by stanbon(75887)

Suppose N=x^+2 models the number ofcases of infection, in millions, of a disease x years from now. Approximately how many years will there be 7 million cases.
----------------
7x10^6 = x^2
x = (sqrt7)(10^3)
x = 2.64575*1000
x = 2645.75 years
====================
Cheers,
Stan H.